OpenStudy (anonymous):

In a property liability case, a company can settle out of court for a loss of $350,000 or go to trial losing $700,000 if found guilty, or lose nothing if found not guilty. lawyers estimate the probability of a not-guilty verdict to be .8 How would I find the expected value of the amount the company can lose by taking the case to court?

OpenStudy (primeralph):

Quite interesting. I have a purely mathematical solution, but I don't wan to use that in an everyday affair like this. @Hero @Zarkon @phi

OpenStudy (anonymous):

Answer is 140,000

hero (hero):

@primeralph, don't hold yourself back.

OpenStudy (primeralph):

@Hero But the bed; it's holding me back.

OpenStudy (phi):

see https://en.wikipedia.org/wiki/Expected_value#Discrete_random_variable.2C_finite_case The expected value of x = amount the company can lose by taking the case to court? \[ E[x]= x_1 \cdot p_1 + x_2 \cdot p_2 \] you have two outcomes: you lose the case (x= 700,000) or you win (x= 0) They say " probability of a not-guilty verdict = .8 " (winning the case) prob of losing the case is 1 - Pr(winning) = 0.2 so the expected loss is \[ E[x]= 0 \cdot 0.8 + 700000 \cdot 0.2 = 140000 \]